OpenStudy (anonymous):

Solve for x: x2 + 24x + 90 = 0 show steps

OpenStudy (anonymous):

By using the quadratic formula: \[\frac{-24 \pm \sqrt{24^{2}- 4 \times 1 \times 90} }{ 2 \times 1 }\]

OpenStudy (anonymous):

Can you do it now?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

Err...do you know the quadratic formula?